i need help with this question
4x-y=-9
-12x+3y=27 may you please help me on it thanks

Answer by Alan3354(69443)

i need help with this question
4x-y=-9
-12x+3y=27
===============
The 2nd eqn is -3 times the first
--> dependent, no unique solution